OpenStudy (anonymous):

Put function into standard form: y=2(x-5)^2+31

OpenStudy (anonymous):

(x-5)^2=x^2-10x+25 y=2(x^2-10x+25)+31 y=2x^2-20x+50+31 y=2x^2-20x+81

OpenStudy (anonymous):

Thank you @marsss

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

@marsss can you explain the steps please ? <3 sorry for all the trouble.

OpenStudy (anonymous):

\[(x-5)^{2}=x ^{2}-10x+25\] \[y=2(x ^{2}-10x+25)+31=2*x^2-2*10x+2*25+50\] does it make sense?

OpenStudy (anonymous):

Yeah , thank you again lol.

OpenStudy (anonymous):

welcome again (: